OpenStudy (anonymous):

bart bought a camera with a list of $219 from an online store offering 6% discount. he needs to pay $7.50 for shipping. what was his total cost?

OpenStudy (anonymous):

\[94\%*219+7.50=0.94*219+7.50=\text{$\$$213}.36 \] If the discount is 6%, then then bart is obligated to pay the remaining 94% plus shipping.